Combine like terms:
x + 6x2 – 4 – x2 – 3 + 9x + 8

Answer by checkley77(12844)

x + 6x2 – 4 – x2 – 3 + 9x + 8
-----------------------------
6x^2-x^2=5x^2
x+9x=10x
-4-3+8=1
-----------------------------
5x^2+10x+1 answer.